# Variational formulations of transport on combinatorial meshes

### v6

@@ 3.3 Domains for physical quantities
We end this section with a short commentary on the domains used for the physical quantities appearing in the exterior calculus models. To simplify discussion we always assumed that all variables are infinitely smooth. This assumption is too strong condition and has to be relaxed in order to guarantee the existence of solutions. For weak formulations the appropriate domains for spaces of $p$-forms are the Sobolev spaces of weakly differentiable $p$-forms [30]
@@ 4.1 Cell complexes (meshes)
Let $(\mathcal{M},\preceq)$ be a mesh. It has been constructed by the requirement of embedding but intrinsically it is a partial order, which we call the topology of $\mathcal{M}$. (A similar definition for CW complexes, requiring only the partial order, is given in [31].) Henceforth, we say that a construction based on the partial order and the relative orientations defined shortly is a topological concept on a mesh. Such concepts will form the backbone of the discrete theory. Considering an intrinsic metric, such as positive numbers associated with the cells of the mesh, will introduce metric concepts, but the theory will still be intrinsic, e.g., independent of any embedding of the mesh in a manifold.
@@
The previous proposition has a direct consequence that any maximal chain for a cell $a$ will have the same number of elements, equal to the dimension of an embedding of $a$ plus $1$. This leads to the following definition of dimension a cell, $\dim a=p$: $p$ can defined extrinsically as the dimension of its embedding, or intrinsically as $p=l-1$, where $l$ is the number of elements a maximal chain for $a$.
@@ 4.2 Topological operations
Let $\mathcal{M}$ be a mesh of dimension $D$ that has the following property: any $(D-1)$-cell is a hyperface of at most two $D$-cells. The $(D-1)$-cells with $2$ adjacent $D$-cells are called interior cells, while those with $1$ adjacent $D$-cell are called boundary cells.) A compatible orientation on $\mathcal{M}$ is a relative orientation $\varepsilon_{D}$ between $D$-cells and $(D-1)$-cells such that for any interior $c\in C^{D-1}\mathcal{M}$ with adjacent $a,b\in C^{D}\mathcal{M}$,
@@
$[\mathcal{M}]:=\sum_{c\in\mathcal{M}_{D}}c_{\bullet}.$ (4.35)
@@
Let $\mathcal{M}$ be a compatibly oriented finite mesh of dimension $D$ with fundamental class $[\mathcal{M}]$, $\sigma\in C^{D}M$. Denote
@@
Cup products exist for simplicial [32], cubical [33], and polygonal [34] meshes. In all cases they are graded commutative and posses the graded Leibniz rule, but are not associative (even in dimension $1$!). However, for simplices the cup product converges to the smooth wedge product [32, Theorem 5.4] and cochains coming from de Rham maps are almost associative [32, Theorem 5.9]. Cup product can also be defined easily for products of meshes with using the tensor product of non-associative graded commutative differential graded algebras, thus allowing to define it over prisms as well. However, we are unaware of a cup product for a generic polytopal mesh. Moreover, we want to use a notion of topological orthogonality which works well for cubical meshes. For this reason we will use a construction of a quasi-cubical mesh $\mathcal{K}$ (i.e., all its cells have face lattices isomorphic to those of cubes) from a generic mesh $\mathcal{M}$ of simple polytopes, called the Forman subdivision, outlined in detail in [26, Section 2]. Another reason to use the Forman subdivision is to allow for more complex material properties (conductivities), as discussed in [26, Section 4]. In this article’s numerical examples we will use only constant conductivities but the intention is to use the discrete formulations (Sections 3.1 and 3.2) with different conductivities of cells of different dimensions as dictated by practical applications.
@@
The topology of the Forman subdivision $\mathcal{K}$ of a mesh $\mathcal{M}$ is defined as follows: the $p$-cells of $\mathcal{K}$ are the intervals $[a,b]:=\{c\in\mathcal{M}\mid a\preceq c\preceq b\}$, where $a\in\mathcal{M}_{q},\ b\in\mathcal{M}_{q-p},\ q\geq p$. The partial order on $\mathcal{K}$ is the sub-interval relation on $\mathcal{M}$, i.e.,
$[a,b]\preceq_{\mathcal{K}}[c,d]\Leftrightarrow c\preceq_{\mathcal{M}}a\preceq_{\mathcal{M}}b\preceq_{\mathcal{M}}d.$ (4.39)
For the embedding of $\mathcal{K}$ it was proposed in [26] to use barycentric coordinates for nodes and flat $p$-cells, if possible for $p\geq 1$. However, in this article we allow generic embeddings and so the natural embedding of $\mathcal{K}$ will be defined in case by case way. For instance, in curved meshes, as those used in Sections 6.2 and 6.3, it will be more natural to use flat cells in the respective curvilinear parametrisations.
@@
if $b$ and $c$ share a single common 00-cell, $b\preceq a$, and $c\preceq a$. Also, for a $(p+q)$-cell $a$, denote by $\perp_{p,q}a$ the set of all pairs of perpendicular $p$- and $q$-subfaces of $a$.
@@
In [26] we worked on a convex mesh $\mathcal{K}$ and relative orientations were defined by vector space orientations of the embedding of $\mathcal{K}$ in $\mathbb{R}^{D}$. More precisely, the expression there was equivalent to
@@
Let $\mathcal{N}$ be the common node of the cubes $a$, $b$ and $c$ so that $a$ be a cube spanned by basis vectors $e_{1},...,e_{p+q}$ with orientation $s_{a}e_{1}\wedge...\wedge e_{p+q}$, $b$ be spanned by $e_{1},...,e_{p}$ with orientation $s_{b}e_{1}\wedge...\wedge e_{p}$, $c$ be spanned by $e_{p+1},...,e_{p+q}$ with orientation $s_{c}e_{p+1}\wedge...\wedge e_{p+q}$, where $s_{a},s_{b},s_{c}\in\{-1,1\}$. Then, substituting in Equation 4.42 leads to $\varepsilon^{\perp}(a,b,c)=s_{a}s_{b}s_{c}$. If $b$ is the 00-cell $\mathcal{N}$ ($s_{b}=1$), then $a=c$ and $s_{A}=s_{c}$, and hence $\varepsilon^{\perp}(a,b,c)=1$. Analogously, if $c$ is the 00-cell $\mathcal{N}$, then $\varepsilon^{\perp}(a,b,c)=1$.
@@
[…]
@@
The above proposition was proved for cubes in [33, Theorem 3.2.3] using a lengthy argument. The same proof can be used there since both identities have to be checked on single quasi-cubical cells (we first map cells of $\mathcal{K}$ to cubes). However, a straightforward proof can be done using the fact that cubical meshes with cup products can be defined as tensor products of graded commutative differential graded algebras (the $1$D cubes (segments) together with their cup products). The base case (a segment) is trivially checked by brute force.
@@ 4.3 Metric operations
Let $(\mathcal{K},\mu)$ be a quasi-cubical Riemannian mesh. The discrete inner product on $\mathcal{K}$ is a family of bilinear forms on $C^{p}\mathcal{K}$ (for $p=0,...,D$) such that basis cochains form an orthogonal basis, and for any $c\in\mathcal{K}_{p}$,
@@
Note that in convex meshes any orthogonal cells $b$ and $c$ have a unique common $D$-cell $a$. Hence, the above formula becomes [35, Equation 2.90]
@@
A proof for a similar formulation on a convex mesh is derived at [26, Equation 16]. Note that in [26, 35, Equation 16] a different convention for Hodge star is used but this one coincides with the standard one, used in exterior algebras.
@@
Let $\mathcal{K}$ be a compatibly oriented quasi-cubical Riemannian mesh, $p\in\{1,...,D\}$, $\sigma\in C^{p}_{0}\mathcal{K},b\in{\rm interior}(\mathcal{K}_{p-1})$. Then [26, Equation 14]
@@
For instance, all topological operators except the wedge/cup products correspond exactly. The accuracy of the correspondence between the wedge product and the cup product on a simplicial mesh is discussed in detail in [32]; one of the crucial differences is the non-associativity of the cup product. For metric operators, all relationships are approximate, which is inescapable. However, our definitions are based on the following principles:
@@
Manifold $M$ $1$ Exact Mesh $\mathcal{M}$ $1$
Submanifold $V$ $1$ Exact Cell $a$ $1$
@@
Boundary $\partial$ $1$ Exact Boundary operator $\partial$ $1$
Exterior derivative $d$ $1$ Exact Coboundary operator $\delta$ $1$
Integration $\int_{M}\omega$ $1$ Exact Discrete integration $\sigma[\mathcal{M}]$ $1$
Trace $\mathop{\rm tr}\nolimits\omega$ $1$ Exact Discrete trace $\mathop{\rm tr}\nolimits\sigma$ $1$
Wedge product $\wedge$ $1$ Approximate Cup product $\smile$ $1$
@@ 5 Combinatorial mesh calculus variational formulations
Discussion 5.1.
The exterior calculus formulations in Section 3 will be translated almost literally into CMC formulations. However, a little bit of care should be taken in light of the fact that the discrete Hodge star operators are not invertible. Hence, in principle, the discrete primal strong, primal weak and mixed weak formulations (both transient and steady-state) are not equivalent to one another. Moreover, their continuous derivations do not translate literally in the discrete setting. However, our numerical tests show good results, and so we will stick to the literal translations.
@@
Quantity Symbol Domain Dimension Pseudo-object? In continuum
Initial potential $u_{0}$ $C^{0}\mathcal{K}$ $\mathsf{Y}$ No eq. 3.36
Internal production rate $f$ $\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.8
Prescribed flow rate $g_{\mathbf{N}}$ $\mathcal{C}^{\infty}(I,C^{D-1}\Gamma_{\mathbf{N}})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.14
Prescribed potential $g_{\mathbf{D}}$ $\mathcal{C}^{\infty}(I,C^{0}\Gamma_{\mathbf{D}})$ $\mathsf{Y}$ No eq. 3.18
Volumetric flow rate $v$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{L}^{D}\mathsf{T}^{-1}$ Yes eq. 3.33
Capacity $\pi$ $\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}$ No eq. 3.21
Dual capacity $\widetilde{\pi}$ $\mathcal{C}^{\infty}(I,C^{0}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{0}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}$ No eq. 3.20
Conductivity $\kappa$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}$ No eq. 3.27
Dual conductivity $\widetilde{\kappa}$ $\mathcal{C}^{\infty}(I,C^{1}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{1}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}$ No eq. 3.30
@@
Amount $Q$ $\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})$ $\mathsf{X}$ Yes eq. 3.3
Flow rate $q$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.5
Diffusive flow rate $q_{\mathcal{D}}$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.24
Advective flow rate $q_{\mathcal{A}}$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.25
Potential $u$ $\mathcal{C}^{\infty}(I,C^{0}\mathcal{K})$ $\mathsf{Y}$ No eq. 3.16
Formulation 5.3 (CMC transient model for transport phenomena).
Combining everything described in this section, under the assumptions of 5.2, we arrive at the model given in Table 8.
@@
$\frac{\partial Q}{\partial t}$ $=f-\delta_{D-1}q$ $\mathsf{X}\mathsf{T}^{-1}$ Conservation law eq. 3.13
$\frac{\partial Q}{\partial t}$ $=\star_{0}\widetilde{\pi}\frac{\partial u}{\partial t}=\pi\star_{0}\frac{\partial u}{\partial t}$ $\mathsf{X}\mathsf{T}^{-1}$ Relation between potential and amount eq. 3.23
$q$ $=q_{\mathcal{D}}+q_{\mathcal{A}}$ $\mathsf{X}\mathsf{T}^{-1}$ Flow rate decomposition eq. 3.26
$q_{\mathcal{D}}$ $=\kappa d^{\star}_{D}\star_{0}u$ $\mathsf{X}\mathsf{T}^{-1}$ Constitutive law eq. 3.29
$q_{\mathcal{A}}$ $=(\star_{D}Q)\smile v$ $\mathsf{X}\mathsf{T}^{-1}$ Advective flow rate formula eq. 3.35
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D},0}}(u)$ $=g_{\mathbf{D}}$ $\mathsf{Y}$ Dirichlet boundary condition eq. 3.19
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N},D-1}}(q)$ $=g_{\mathbf{N}}$ $\mathsf{X}\mathsf{T}^{-1}$ Neumann boundary condition eq. 3.15
$u(t_{0})$ $=u_{0}$ $\mathsf{Y}$ Initial condition eq. 3.37

### v7

@@ 3.3 Domains for physical quantities
We end this section with a short commentary on the domains used for the physical quantities appearing in the exterior calculus models. To simplify discussion we always assumed that all variables are infinitely smooth. This assumption is too strong condition and has to be relaxed in order to guarantee the existence of solutions. For weak formulations the appropriate domains for spaces of $p$-forms are the Sobolev spaces of weakly differentiable $p$-forms arnold2018finite
@@ 4.1 Cell complexes (meshes)
Let $(\mathcal{M},\preceq)$ be a mesh. It has been constructed by the requirement of embedding but intrinsically it is a partial order, which we call the topology of $\mathcal{M}$. (A similar definition for CW complexes, requiring only the partial order, is given in bjorner1984posets .) Henceforth, we say that a construction based on the partial order and the relative orientations defined shortly is a topological concept on a mesh. Such concepts will form the backbone of the discrete theory. Considering an intrinsic metric, such as positive numbers associated with the cells of the mesh, will introduce metric concepts, but the theory will still be intrinsic, e.g., independent of any embedding of the mesh in a manifold.
@@
The previous proposition has a direct consequence that any maximal chain for a cell $a$ will have the same number of elements, equal to the dimension of an embedding of $a$ plus $1$. This leads to the following definition of dimension a cell, $\dim a=p$: $p$ can defined extrinsically as the dimension of its embedding, or intrinsically as p=l1p=l-1, where $l$ is the number of elements a maximal chain for $a$.
@@ 4.2 Topological operations
Let $\mathcal{M}$ be a mesh of dimension $D$ that has the following property: any (D1)(D-1)-cell is a hyperface of at most two $D$-cells. The (D1)(D-1)-cells with $2$ adjacent $D$-cells are called interior cells, while those with $1$ adjacent $D$-cell are called boundary cells.) A compatible orientation on $\mathcal{M}$ is a relative orientation $\varepsilon_{D}$ between $D$-cells and (D1)(D-1)-cells such that for any interior $c\in C^{D-1}\mathcal{M}$ with adjacent $a,b\in C^{D}\mathcal{M}$,
@@
$\mathcal{M}]:=\sum_{c\in\mathcal{M}_{D}}c_{\bullet}.$ (4.35)
@@
Let $\mathcal{M}$ be a compatibly oriented finite mesh of dimension $D$ with fundamental class $\mathcal{M}$], $\sigma\in C^{D}M$. Denote
@@
Cup products exist for simplicial wilson2007cochain , cubical arnold2012discrete , and polygonal ptackova2017discrete meshes. In all cases they are graded commutative and posses the graded Leibniz rule, but are not associative (even in dimension 11!). However, for simplices the cup product converges to the smooth wedge product [wilson2007cochain, , Theorem 5.4] and cochains coming from de Rham maps are almost associative [wilson2007cochain, , Theorem 5.9]. Cup product can also be defined easily for products of meshes with using the tensor product of non-associative graded commutative differential graded algebras, thus allowing to define it over prisms as well. However, we are unaware of a cup product for a generic polytopal mesh. Moreover, we want to use a notion of topological orthogonality which works well for cubical meshes. For this reason we will use a construction of a quasi-cubical mesh $\mathcal{K}$ (i.e., all its cells have face lattices isomorphic to those of cubes) from a generic mesh $\mathcal{M}$ of simple polytopes, called the Forman subdivision, outlined in detail in [berbatov2022diffusion, , Section 2]. Another reason to use the Forman subdivision is to allow for more complex material properties (conductivities), as discussed in [berbatov2022diffusion, , Section 4]. In this article’s numerical examples we will use only constant conductivities but the intention is to use the discrete formulations (Sections 3.1 and 3.2) with different conductivities of cells of different dimensions as dictated by practical applications.
@@
The topology of the Forman subdivision $\mathcal{K}$ of a mesh $\mathcal{M}$ is defined as follows: the $p$-cells of $\mathcal{K}$ are the intervals $a,b]:=\{c\in\mathcal{M}\mid a\preceq c\preceq b\}$, where $a\in\mathcal{M}_{q},\ b\in\mathcal{M}_{q-p},\ q\geq p$. The partial order on $\mathcal{K}$ is the sub-interval relation on $\mathcal{M}$, i.e.,
$a,b]\preceq_{\mathcal{K}}[c,d]\Leftrightarrow c\preceq_{\mathcal{M}}a\preceq_{\mathcal{M}}b\preceq_{\mathcal{M}}d.$ (4.39)
For the embedding of $\mathcal{K}$ it was proposed in berbatov2022diffusion to use barycentric coordinates for nodes and flat $p$-cells, if possible for $p\geq 1$. However, in this article we allow generic embeddings and so the natural embedding of $\mathcal{K}$ will be defined in case by case way. For instance, in curved meshes, as those used in Sections 6.2 and 6.3, it will be more natural to use flat cells in the respective curvilinear parametrisations.
@@
if $b$ and $c$ share a single common 0-cell, $b\preceq a$, and $c\preceq a$. Also, for a $(p+q)$-cell $a$, denote by $\perp_{p,q}a$ the set of all pairs of perpendicular $p$- and $q$-subfaces of $a$.
@@
In berbatov2022diffusion we worked on a convex mesh $\mathcal{K}$ and relative orientations were defined by vector space orientations of the embedding of $\mathcal{K}$ in $\mathbb{R}^{D}$. More precisely, the expression there was equivalent to
@@
Let $\mathcal{N}$ be the common node of the cubes $a$, $b$ and $c$ so that $a$ be a cube spanned by basis vectors $e_{1},...,e_{p+q}$ with orientation $s_{a}e_{1}\wedge...\wedge e_{p+q}$, $b$ be spanned by $e_{1},...,e_{p}$ with orientation $s_{b}e_{1}\wedge...\wedge e_{p}$, $c$ be spanned by $e_{p+1},...,e_{p+q}$ with orientation $s_{c}e_{p+1}\wedge...\wedge e_{p+q}$, where $s_{a},s_{b},s_{c}\in\{-1,1\}$. Then, substituting in Equation 4.42 leads to $\varepsilon^{\perp}(a,b,c)=s_{a}s_{b}s_{c}$. If $b$ is the 0-cell $\mathcal{N}$ ($s_{b}=1$), then $a=c$ and $s_{A}=s_{c}$, and hence $\varepsilon^{\perp}(a,b,c)=1$. Analogously, if $c$ is the 0-cell $\mathcal{N}$, then $\varepsilon^{\perp}(a,b,c)=1$.
@@
$\smile\colon C^{\bullet}\mathcal{K}\times C^{\bullet}\mathcal{K}\to C^{\bullet}\mathcal{K},\ \smile_{p,q}\colon C^{p}\mathcal{K}\times C^{q}\mathcal{K}\to C^{p+q}\mathcal{K},\ 0\leq p,q\leq D,\ p+q\leq D$ (4.50)
@@
The above proposition was proved for cubes in [arnold2012discrete, , Theorem 3.2.3] using a lengthy argument. The same proof can be used there since both identities have to be checked on single quasi-cubical cells (we first map cells of $\mathcal{K}$ to cubes). However, a straightforward proof can be done using the fact that cubical meshes with cup products can be defined as tensor products of graded commutative differential graded algebras (the 11D cubes (segments) together with their cup products). The base case (a segment) is trivially checked by brute force.
@@ 4.3 Metric operations
Let $(\mathcal{K},\mu)$ be a quasi-cubical Riemannian mesh. The discrete inner product on $\mathcal{K}$ is a family of bilinear forms on $C^{p}\mathcal{K}$ (for p=0,,Dp=0,...,D) such that basis cochains form an orthogonal basis, and for any $c\in\mathcal{K}_{p}$,
@@
Note that in convex meshes any orthogonal cells $b$ and $c$ have a unique common $D$-cell $a$. Hence, the above formula becomes [berbatov2023discrete, , Equation 2.90]:
@@
A proof for a similar formulation on a convex mesh is derived at [berbatov2022diffusion, , Equation 16]. Note that in berbatov2022diffusion , berbatov2023discrete a different convention for Hodge star is used but this one coincides with the standard one, used in exterior algebras.
@@
Let $\mathcal{K}$ be a compatibly oriented quasi-cubical Riemannian mesh, $p\in\{1,...,D\}$, $\sigma\in C^{p}_{0}\mathcal{K},b\in{\rm interior}(\mathcal{K}_{p-1})$. Then [berbatov2022diffusion, , Equation 14]
@@
For instance, all topological operators except the wedge/cup products correspond exactly. The accuracy of the correspondence between the wedge product and the cup product on a simplicial mesh is discussed in detail in wilson2007cochain ; one of the crucial differences is the non-associativity of the cup product. For metric operators, all relationships are approximate, which is inescapable. However, our definitions are based on the following principles:
@@
Manifold $M$ 11 Exact Mesh $\mathcal{M}$ 11
Submanifold $V$ 11 Exact Cell $a$ 11
@@
Boundary $\partial$ 11 Exact Boundary operator $\partial$ 11
Exterior derivative $d$ 11 Exact Coboundary operator $\delta$ 11
Integration $\int_{M}\omega$ 11 Exact Discrete integration $\sigma[\mathcal{M}]$ 11
Trace $\mathop{\rm tr}\nolimits\omega$ 11 Exact Discrete trace $\mathop{\rm tr}\nolimits\sigma$ 11
Wedge product $\wedge$ 11 Approximate Cup product $\smile$ 11
@@ 5 Combinatorial mesh calculus variational formulations
Remark 5.1.
The exterior calculus formulations in Section 3 can be translated almost literally into CMC formulations. However, we note that the discrete Hodge star operators are not invertible. Therefore, in principle, the discrete strong, primal weak and mixed weak formulations (both transient and steady-state) are not equivalent to one another. Their continuous derivations do not translate literally in the discrete setting. Because our numerical tests show good results, we use literal translations in this work.
@@
Quantity Symbol Domain Dimension Pseudo-object? Continuum analogue
Initial potential $u_{0}$ $C^{0}\mathcal{K}$ $\mathsf{Y}$ No Equation 3.36
Internal production rate $f$ $\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.8
Prescribed flow rate $g_{\mathbf{N}}$ $\mathcal{C}^{\infty}(I,C^{D-1}\Gamma_{\mathbf{N}})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.14
Prescribed potential $g_{\mathbf{D}}$ $\mathcal{C}^{\infty}(I,C^{0}\Gamma_{\mathbf{D}})$ $\mathsf{Y}$ No Equation 3.18
Volumetric flow rate $v$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{L}^{D}\mathsf{T}^{-1}$ Yes Equation 3.33
Capacity $\pi$ $\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}$ No Equation 3.21
Dual capacity $\widetilde{\pi}$ $\mathcal{C}^{\infty}(I,C^{0}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{0}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}$ No Equation 3.20
Conductivity $\kappa$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}$ No Equation 3.27
Dual conductivity $\widetilde{\kappa}$ $\mathcal{C}^{\infty}(I,C^{1}\mathcal{K})\to\mathcal{C}^{\infty}(I,C^{1}\mathcal{K})$ $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}$ No Equation 3.30
@@
Amount $Q$ $\mathcal{C}^{\infty}(I,C^{D}\mathcal{K})$ $\mathsf{X}$ Yes Equation 3.3
Flow rate $q$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.5
Diffusive flow rate $q_{\mathcal{D}}$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.24
Advective flow rate $q_{\mathcal{A}}$ $\mathcal{C}^{\infty}(I,C^{D-1}\mathcal{K})$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.25
Potential $u$ $\mathcal{C}^{\infty}(I,C^{0}\mathcal{K})$ $\mathsf{Y}$ No Equation 3.16
Remark 5.3.
The input physical quantities ($u_{0}$, $f$, $g_{\mathbf{N}}$, $g_{\mathbf{D}}$, and $v$) can be provided either directly in discrete form or as discretised versions of their continuous counterparts via the de Rham map – see 6.2 for further details. Material properties are assigned to the cells of the complex: capacity is associated with $3$-cells (and its dual with 0-cells), while conductivity is associated with (D1)(D-1)-cells (and its dual with $1$-cells). For example, for a $1$-cell $c$, we define $(\tilde{\kappa}q)(c)=\lambda_{c}\,q(c)$, where $\lambda_{c}>0$ is a material-specific coefficient. Analogous relations hold for the operators $\pi$, $\kappa$, and $\tilde{\pi}$.
Our approach is designed to predict or derive macroscopic (emergent) properties of the complex from given local, cell-wise properties. This enables the analysis of heterogeneous structures with arbitrary geometric and topological complexity. The inverse problem – determining local properties that give rise to prescribed macroscopic behaviour – is a compelling and important challenge, especially in the context of materials design with target functionalities. While addressing this inverse problem is beyond the scope of the present work, it could be pursued via various machine learning strategies.
In this paper, we consider the simpler case of homogeneous and isotropic materials with prescribed macroscopic conductivity and capacity. In such cases, the inverse problem is trivial: all cells associated with conductivity or capacity are assigned the corresponding macroscopic values. This setup enables direct comparison with classical continuum models in our simulation examples.
It is important to note that, since the discrete Hodge star operators are not invertible, material coefficients and their dual counterparts are not interchangeable. In the following two points, we explain how conductivities and capacities are defined in the discrete setting.
Discussion 5.4 (Specifying discrete conductivities).
Following berbatov2022diffusion , we define conductivity on $1$-cells – this corresponds to the dual conductivity $\tilde{\kappa}$ in this work. The $1$-cells in $\mathcal{K}$ correspond to intervals $c$ in the physical complex $\mathcal{M}$, where $\dim b=\dim a+1$ for $[a,b]=c$.
Conductivities are specified as positive scalar values on non-nodal cells of $\mathcal{M}$. Each $1$-cell $[a,b]$ in $\mathcal{K}$ is assigned the conductivity value of the target cell $b$. The corresponding primal conductivity $\kappa$ is represented as a diagonal matrix, determined via a least-squares fit to satisfy either $\star_{1}\,\tilde{\kappa}=\kappa\,\star_{1}$ or $\star_{D-1}\,\kappa=\tilde{\kappa}\,\star_{D-1}$.
Discussion 5.5 (Specifying discrete capacities).
In parallel with the conductivity specification, we begin by assigning the dual capacity $\tilde{\pi}$. Note that 0-cells in $\mathcal{K}$ correspond to all cells in $\mathcal{M}$. Capacities are given as positive scalar values on $p$-cells of $\mathcal{M}$ for $p>0$ only. For any 0-cell $[a,a]$ in $\mathcal{K}$, we assign the capacity of $a$ if $a$ is not a 0-cell of $\mathcal{M}$, and zero otherwise. The corresponding primal capacity $\pi$ is again represented as a diagonal matrix, determined via a least-squares fit to satisfy either $\star_{0}\,\tilde{\pi}=\pi\,\star_{0}$ or $\star_{D}\,\pi=\tilde{\pi}\,\star_{D}$.
Formulation 5.6 (CMC transient model for transport phenomena).
By translating Table 4 into the language of combinatorial meshes, under the assumptions of 5.2, we arrive at the model given in Table 8.
@@
$\frac{\partial Q}{\partial t}$ $f-\delta_{D-1}q$ $\mathsf{X}\mathsf{T}^{-1}$ Conservation law Equation 3.13
$\frac{\partial Q}{\partial t}$ $\star_{0}\widetilde{\pi}\frac{\partial u}{\partial t}=\pi\star_{0}\frac{\partial u}{\partial t}$ $\mathsf{X}\mathsf{T}^{-1}$ Relation between potential and amount Equation 3.23
$q$ $q_{\mathcal{D}}+q_{\mathcal{A}}$ $\mathsf{X}\mathsf{T}^{-1}$ Flow rate decomposition Equation 3.26
$q_{\mathcal{D}}$ $\kappa d^{\star}_{D}\star_{0}u$ $\mathsf{X}\mathsf{T}^{-1}$ Constitutive law Equation 3.29
$q_{\mathcal{A}}$ $(\star_{D}Q)\smile v$ $\mathsf{X}\mathsf{T}^{-1}$ Advective flow rate formula Equation 3.35
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D},0}}(u)$ $g_{\mathbf{D}}$ $\mathsf{Y}$ Dirichlet boundary condition Equation 3.19
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N},D-1}}(q)$ $g_{\mathbf{N}}$ $\mathsf{X}\mathsf{T}^{-1}$ Neumann boundary condition Equation 3.15
$u(t_{0})$ $u_{0}$ $\mathsf{Y}$ Initial condition Equation 3.37